# The entries of the Sinkhorn limit of an $m\times n$ matrix

### v1

@@ 4. Rectangular matrices and combinatorial structure
The coefficients $c_{S}(n)$ for $n\in\{2,3,4,5\}$ that we have computed so far are not enough to guess the general formula for $c_{S}(n)$. In this section, we expand the scope to include matrices that are not necessarily square. This allows us to compute enough coefficients to identify formulas in special cases. These formulas reveal the relevant combinatorial structure on subsets $S$, and this allows us to piece together all but one detail of the general formula for the coefficients, resulting in Conjecture 2.
@@
Conjecture 9 generalizes to $m\times n$ matrices as follows. We extend the coefficients $c_{S}(n)$ in the previous section to coefficients $c_{S}(m,n)$, where $c_{S}(n,n)=c_{S}(n)$. For each $m$ and $n$, we scale the coefficients so that $c_{\{\}}(m,n)=1$. We have $\lvert{D(m,n)}\rvert=\sum_{k=0}^{\min(m,n)-1}\binom{m-1}{k}\binom{n-1}{k}=\binom{m+n-2}{m-1}$.
@@
Let $m\geq 1$ and $n\geq 1$. There exist rational numbers $c_{S}(m,n)$, indexed by subsets $S\subseteq D(m,n)$, such that, for every positive $m\times n$ matrix $A$, the top left entry $x$ of $\operatorname{Sink}(A)$ satisfies
@@
In particular, $x$ has degree at most $\binom{m+n-2}{m-1}$.
To interpolate values of $c_{S}(m,n)$, we use the method described in Section 3. Overall, we used $1.5$ years of CPU time to iteratively scale matrices of various sizes and recognize $102000$ algebraic numbers. An additional month of CPU time was spent setting up and solving systems of linear equations in the coefficients $c_{S}(m,n)$, which resulted in the identification of $63000$ rational coefficients (and an additional $56000$ coefficients parameterized by free variables).
@@
There is an asymmetry in our definition of $\operatorname{Sink}(A)$ for non-square matrices, since $\operatorname{Sink}(A)$ has row sums $1$ and column sums $\frac{m}{n}$. However, we expect some symmetry in the coefficients $c_{S}(m,n)$ since $m\operatorname{Sink}(A^{\textnormal{{T}}})=n\operatorname{Sink}(A)^{\textnormal{{T}}}$. We can obtain this symmetry by considering $\frac{1}{m}\operatorname{Sink}(A)$, which has row sums $\frac{1}{m}$ and column sums $\frac{1}{n}$. The form of the polynomial for the top left entry $y$ of $\frac{1}{m}\operatorname{Sink}(A)$ can be obtained from Conjecture 12 by substituting $x=my$. Therefore, the coefficients $c_{S}(m,n)$ should satisfy $m^{\lvert{S}\rvert}c_{S}(m,n)=n^{\lvert{S}\rvert}c_{S^{\textnormal{{T}}}}(n,m)$, where $S^{\textnormal{{T}}}$ is defined by
@@
For $S=\begin{smallmatrix}\{2\}\\ \{2\}\end{smallmatrix}$, the data suggests $mc_{S}(m,n)=m+n-mn$ (for all $m\geq 2$ and $n\geq 2$). For $S=\begin{smallmatrix}\{2\}\\ \{3\}\end{smallmatrix}$, the data suggests the same formula $mc_{S}(m,n)=m+n-mn)$; indeed this is expected, since $\begin{smallmatrix}\{2\}\\ \{2\}\end{smallmatrix}\equiv\begin{smallmatrix}\{2\}\\ \{3\}\end{smallmatrix}$. Since each subset $\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}$ with $\lvert{R}\rvert=i$ is equivalent to $S=\begin{smallmatrix}\{2,3,\dots,i+1\}\\ \{2,3,\dots,i+1\}\end{smallmatrix}$, it suffices to consider the latter. For $S=\begin{smallmatrix}\{2,3\}\\ \{2,3\}\end{smallmatrix}$, the data suggests $mc_{S}(m,n)=2m+2n-mn$. For $S=\begin{smallmatrix}\{2,3,4\}\\ \{2,3,4\}\end{smallmatrix}$, it suggests $mc_{S}(m,n)=3m+3n-mn$.
@@
Let $S=\begin{smallmatrix}\{2\}&\{2\}\\ \{2\}&\{3\}\end{smallmatrix}$. The two minor specifications in $S$ involve the same row but different columns. Therefore there is no subset $T$ in the equivalence class of $S$ such that $T^{\textnormal{{T}}}=T$, so we do not expect $m^{2}c_{S}(m,n)$ to be a symmetric function of $m$ and $n$. Here are several of its values (where the bottom right entry is not included because we have no data for $6\times 6$ matrices):
@@
    • $R_{1}=R_{2}\setminus\{i\}$ and $C_{1}=C_{2}\setminus\{j\}$ for some $i\in R_{2}$ and $j\in C_{2}$, or
    • $R_{2}=R_{1}\setminus\{i\}$ and $C_{2}=C_{1}\setminus\{j\}$ for some $i\in R_{1}$ and $j\in C_{1}$.
@@
    • $R_{1}=R_{2}$ and $C_{1}\setminus\{i\}=C_{2}\setminus\{j\}$ for some $i\in C_{1}$ and $j\in C_{2}$ such that $i\neq j$, or
    • $C_{1}=C_{2}$ and $R_{1}\setminus\{i\}=R_{2}\setminus\{j\}$ for some $i\in R_{1}$ and $j\in R_{2}$ such that $i\neq j$.
@@
Example 19.
Let $S=\begin{smallmatrix}\{2\}&\{2,3\}\\ \{2\}&\{2,3\}\end{smallmatrix}$. The two minor specifications ($(\{2\},\{2\})$ and ($(\{2,3\},\{2,3\})$ form a type-1 link. The values of $m^{2}c_{S}(m,n)$ suggest $m^{2}c_{S}(m,n)=(m+2n-mn)(2m+n-mn)$.
Example 20.
Let $S=\begin{smallmatrix}\{2,3\}&\{2,4\}\\ \{2,3\}&\{2,3\}\end{smallmatrix}$, which consists of a type-2 link. The data suggests $m^{2}c_{S}(m,n)=(2m+n-mn)(2m+3n-mn)$.
Example 21.
Let $S=\begin{smallmatrix}\{2\}&\{2,3,4\}\\ \{2\}&\{2,3,4\}\end{smallmatrix}$, whose two elements are not linked. The data suggests $m^{2}c_{S}(m,n)=(m+n-mn)(3m+3n-mn)$.
@@
Conjectures 16 and 22 imply that, if $\lvert{S}\rvert\leq 2$, then the polynomial $m^{\lvert{S}\rvert}c_{S}(m,n)$ is a product of factors that are linear in $m$ and linear in $n$. When we consider subsets $S$ with $\lvert{S}\rvert\geq 3$, we find additional polynomials with this property. In particular, when $S$ has no linked pairs, there seems to be a simple description of $m^{\lvert{S}\rvert}c_{S}(m,n)$ as follows.
@@
This conjecture suggests more generally that the structure of the polynomial $m^{k}c_{S}(m,n)$ is determined by the connected components of linked pairs in $S$ and, moreover, that the contributions of the components are independent of each other. This is the content of Conjecture 25 below. We make this precise by defining the following graph.
@@
Let $S=\begin{smallmatrix}\{\}&\{2\}&\{2,3\}&\{2,4\}\\ \{\}&\{2\}&\{3,4\}&\{3,4\}\end{smallmatrix}$. The first two elements of $S$ form a type-1 link, and the last two form a type-2 link. These are the only links, so $G_{S}$ is the graph on $4$ vertices with two non-adjacent edges. The two connected components are equivalent to the subsets in Examples 17 and 20, respectively. The values we computed of $m^{4}c_{S}(m,n)$ are as follows.
@@
This is consistent with $m^{4}c_{S}(m,n)=(m-mn)(n-mn)(2m+n-mn)(2m+3n-mn)$, which is the product of the formulas in Examples 17 and 20.
@@
Let $S=\begin{smallmatrix}\{2\}&\{3\}&\{2,3\}\\ \{2\}&\{3\}&\{2,3\}\end{smallmatrix}$. The minor specifications ($(\{2\},\{2\})$ and ($(\{3\},\{3\})$ are not linked, but each of the other two pairs is. Therefore $G_{S}$ consists of a single connected component. Here are several values of $m^{3}c_{S}(m,n)$:
@@
Without the value for $6\times 6$ matrices, we cannot interpolate a cubic polynomial in $m$ and $n$. However, by interpolating cubic polynomials in $n$ for the first three rows, we see that $m+n-mn$ is likely a factor. Dividing each value in the table by this factor, we then interpolate a quadratic polynomial to obtain $m^{3}c_{S}(m,n)=(m+n-mn)\left(2m^{2}+2n^{2}+6mn-3m^{2}n-3mn^{2}+m^{2}n^{2}\right)$. This quadratic factor is irreducible.
An obvious question is whether there is a better way to write such polynomials, so that we can see the general structure. In fact there is, using determinants. Since the determinant of a block diagonal matrix is the product of the determinants of the blocks, determinant formulas are good candidates for functions that decompose as products over connected components. In this direction, we next rewrite Conjecture 22 using determinant formulas; these are likely more natural than the factorizations in Conjecture 22.
@@
    $m^{2}c_{S}(m,n)=\det\begin{bmatrix}\lvert{R_{1}}\rvert(m+n)-mn&-m\\ -m&\lvert{R_{2}}\rvert(m+n)-mn\end{bmatrix}.$
@@
    $m^{2}c_{S}(m,n)=\det\begin{bmatrix}\lvert{R_{1}}\rvert(m+n)-mn&-n\\ -n&\lvert{R_{2}}\rvert(m+n)-mn\end{bmatrix}.$
@@
Along with Conjecture 25, this suggests a determinant formula for $m^{\lvert{S}\rvert}c_{S}(m,n)$ for an arbitrary subset $S$. Since the matrix in this determinant formula resembles an adjacency matrix, we introduce the following notation.
@@
For each $S=\begin{smallmatrix}R_{1}&R_{2}&\cdots&R_{k}\vphantom{\{}\\ C_{1}&C_{2}&\cdots&C_{k}\vphantom{\{}\end{smallmatrix}$, define $\operatorname{adj}_{S}(m,n)$ to be the $k\times k$ matrix with the property that, for all $i,j$ satisfying $1\leq i<j\leq k$, the $2\times 2$ submatrix $(\operatorname{adj}_{S}(m,n))_{\{i,j\},\{i,j\}}$ is the matrix in Conjecture 27 for $(R_{i},C_{i})$ and $(R_{j},C_{j})$. In particular, the $i$th diagonal entry of $\operatorname{adj}_{S}(m,n)$ is $\lvert{R_{i}}\rvert(m+n)-mn$, and the off-diagonal entries are elements of $\{-m,-n,0,m\}$.
@@
$\operatorname{adj}_{S}(m,n)=\begin{bmatrix}m+n-mn&0&m\\ 0&m+n-mn&m\\ -n&-n&2m+2n-mn\end{bmatrix}.$
Indeed, $\det\operatorname{adj}_{S}(m,n)$ is equivalent to the formula in Example 26 for $m^{3}c_{S}(m,n)$.
Unfortunately, this construction doesn’t always work.
@@
Let $S=\begin{smallmatrix}\{2,3\}&\{2,3\}&\{2,3\}\\ \{2,3\}&\{2,4\}&\{2,5\}\end{smallmatrix}$. Each pair of elements in $S$ forms a type-2 link, and the common column is the same for all three links. We have
$\operatorname{adj}_{S}(m,n)=\begin{bmatrix}2m+2n-mn&-m&-m\\ -m&2m+2n-mn&-m\\ -m&-m&2m+2n-mn\end{bmatrix}.$
The formula $\det\operatorname{adj}_{S}(m,n)$ does not produce the following values we computed for $m^{3}c_{S}(m,n)$.
$\begin{array}[]{r|rrrrrrr}&n=5&6&7&8&9&10&11\\ \hline\cr m=3&28&54&80&100&108&98&64\\ 4&216&256&200&0\\ 5&500&338\\ 6&784\end{array}$
However, if we alter the signs of the off-diagonal terms, we can get a determinant formula that produces these values, namely
$m^{3}c_{S}(m,n)=\det\begin{bmatrix}2m+2n-mn&m&m\\ m&2m+2n-mn&m\\ m&m&2m+2n-mn\end{bmatrix}.$
We conjecture below that the signs of the off-diagonal terms in $\operatorname{adj}_{S}(m,n)$ can always be altered in such a way, independent of $m$ and $n$, that its determinant gives the value of $m^{\lvert{S}\rvert}c_{S}(m,n)$. Exactly how to alter them is the detail we have not been able to determine. This alteration is not unique, since if we pick a set $V\subseteq\{1,2,\dots,k\}$ of indices and negate the rows and columns indexed by $V$ then the determinant does not change. An equivalence class of correct sign alterations is generated in this way, all of which lead to the value of $m^{\lvert{S}\rvert}c_{S}(m,n)$. This equivalence class seems to depend only on the link structure of $S$ and not on the sizes of its elements.
@@
Let $S=\begin{smallmatrix}\{2,3,4\}&\{2,3,4\}&\{2,3,4\}\\ \{2,3,4\}&\{2,3,5\}&\{2,3,6\}\end{smallmatrix}$. The elements of $S$ are related to each other in the same way as in the previous example; each pair forms a type-2 link, and the common columns are the same for all three links. Using the same signs as in the previous example, the determinant formula
$m^{3}c_{S}(m,n)=\det\begin{bmatrix}3m+3n-mn&m&m\\ m&3m+3n-mn&m\\ m&m&3m+3n-mn\end{bmatrix}$
agrees with the $4$ values we computed. These same signs also work for $S=\begin{smallmatrix}\{2\}&\{2\}&\{2\}\\ \{2\}&\{3\}&\{4\}\end{smallmatrix}$, which also has the same link structure.
Our most general result is therefore Conjecture 2, which appears in Section 1 and which we restate below with more information. We introduce one last bit of notation to enable sign alterations. We also take the opportunity to scale the entries of $\operatorname{adj}_{S}(m,n)$ by $\frac{1}{m}$; this allows us to dispense with the factor $m^{\lvert{S}\rvert}$ that would otherwise appear in the summand in Conjecture 2.
Notation.
Let $S\subseteq D(m,n)$, and let $k=\lvert{S}\rvert$. Let $\sigma\colon\{1,2,\dots,k\}^{2}\to\{-1,1\}$. Define $\operatorname{adj}_{S,\sigma}(m,n)$ to be the $k\times k$ matrix whose $(i,j)$ entry is $\frac{1}{m}\sigma((i,j))$ times the $(i,j)$ entry of $\operatorname{adj}_{S}(m,n)$.
Conjecture 2.
Let $m\geq 1$ and $n\geq 1$. For each $S\subseteq D(m,n)$, there exists a function $\sigma(S)\colon\{1,2,\dots,\lvert{S}\rvert\}^{2}\to\{-1,1\}$ with $\sigma(S)((i,i))=1$ for all $i$ such that, for every positive $m\times n$ matrix $A$, the top left entry $x$ of $\operatorname{Sink}(A)$ satisfies
$\sum_{S\subseteq D(m,n)}\left(\det\operatorname{adj}_{S,\sigma(S)}(m,n)\right)M(S)x^{\lvert{S}\rvert}=0.$
Moreover, the sign alterations $\sigma(S)$ can be chosen so that they
  • satisfy $\sigma(T)=\sigma(S)$ for all $T\equiv S$,
  • are independent of $m$ and $n$, and
  • depend only on the link structure of $S$ and not on the sizes of its elements.
We conclude this section by identifying the correct sign alterations for $2\times n$ matrices for general $n$, leading to an explicit formula for the equation satisfied by $x$. For each exponent $k$, there are at most $2$ equivalence classes of subsets — an equivalence class containing subsets $S$ such that ({},{}$(\{\},\{\})\in S$ and another containing the rest. We consider them separately.

### v2

@@ 4. Rectangular matrices and combinatorial structure
The coefficients $c_{S}(n)$ for $n\in\{2,3,4,5\}$ that we have computed so far are not enough to guess the general formula for $c_{S}(n)$. In this section, we expand the scope to include matrices that are not square. This allows us to compute enough coefficients to identify formulas in special cases. These formulas reveal the relevant combinatorial structure on subsets $S$, and this allows us to piece together a general formula for the coefficients, resulting in Conjecture 2.
@@
Conjecture 9 generalizes to $m\times n$ matrices as follows. We extend the coefficients $c_{S}(n)$ in the previous section to coefficients $c_{S}(m,n)$, where $c_{S}(n,n)=c_{S}(n)$. For each $m$ and $n$, we scale the coefficients so that $c_{\{\}}(m,n)=1$.
@@
Let $m\geq 1$ and $n\geq 1$. There exist rational numbers $c_{S}(m,n)$, indexed by subsets $S\subseteq D(m,n)$, such that, for every positive $m\times n$ matrix $A$, the top-left entry $x$ of $\operatorname{Sink}(A)$ satisfies
@@
To interpolate values of $c_{S}(m,n)$, we use the method described in Section 3. Overall, we spent $1.5$ years of CPU time iteratively scaling matrices of various dimensions and recognizing $102000$ algebraic numbers. An additional month of CPU time was spent setting up and solving systems of linear equations in the coefficients $c_{S}(m,n)$, which resulted in the identification of $63000$ rational coefficients (and an additional $56000$ coefficients parameterized by free variables).
@@
There is an asymmetry in our definition of $\operatorname{Sink}(A)$ for non-square matrices, since $\operatorname{Sink}(A)$ has row sums $1$ and column sums $\frac{m}{n}$. However, we expect some symmetry in the coefficients $c_{S}(m,n)$ since $m\operatorname{Sink}(A^{\textnormal{{T}}})=n\operatorname{Sink}(A)^{\textnormal{{T}}}$. We can obtain this symmetry by considering $\frac{1}{m}\operatorname{Sink}(A)$, which has row sums $\frac{1}{m}$ and column sums $\frac{1}{n}$. The form of the polynomial for the top-left entry $y$ of $\frac{1}{m}\operatorname{Sink}(A)$ can be obtained from Conjecture 14 by substituting $x=my$. Therefore, the coefficients $c_{S}(m,n)$ should satisfy $m^{\lvert{S}\rvert}c_{S}(m,n)=n^{\lvert{S}\rvert}c_{S^{\textnormal{{T}}}}(n,m)$, where $S^{\textnormal{{T}}}$ is defined by
@@
For $S=\begin{smallmatrix}\{2\}\\ \{2\}\end{smallmatrix}$, the data suggests $mc_{S}(m,n)=m+n-mn$ (for all $m\geq 2$ and $n\geq 2$). For $S=\begin{smallmatrix}\{2\}\\ \{3\}\end{smallmatrix}$, the data suggests the same formula $mc_{S}(m,n)=m+n-mn$; indeed this is expected, because $\begin{smallmatrix}\{2\}\\ \{2\}\end{smallmatrix}\equiv\begin{smallmatrix}\{2\}\\ \{3\}\end{smallmatrix}$. Since every size-$1$ subset $\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}$ is equivalent to $S=\begin{smallmatrix}\{2,3,\dots,\lvert{R}\rvert+1\}\\ \{2,3,\dots,\lvert{R}\rvert+1\}\end{smallmatrix}$, it suffices to consider the latter. For $S=\begin{smallmatrix}\{2,3\}\\ \{2,3\}\end{smallmatrix}$, the data suggests $mc_{S}(m,n)=2m+2n-mn$. For $S=\begin{smallmatrix}\{2,3,4\}\\ \{2,3,4\}\end{smallmatrix}$, it suggests $mc_{S}(m,n)=3m+3n-mn$.
@@
Let $S=\begin{smallmatrix}\{2\}&\{2\}\\ \{2\}&\{3\}\end{smallmatrix}$. The two minor specifications in $S$ involve the same row but different columns. Therefore there is no subset $T$ in the equivalence class of $S$ such that $T^{\textnormal{{T}}}=T$, so we do not expect $m^{2}c_{S}(m,n)$ to be a symmetric function of $m$ and $n$. Here are several of its values (where the bottom-right entry is not included because we have no data for $6\times 6$ matrices):
@@
    • $R_{1}=R_{2}\setminus\{s\}$ and $C_{1}=C_{2}\setminus\{t\}$ for some $s\in R_{2}$ and $t\in C_{2}$, or
    • $R_{2}=R_{1}\setminus\{s\}$ and $C_{2}=C_{1}\setminus\{t\}$ for some $s\in R_{1}$ and $t\in C_{1}$.
@@
    • $R_{1}=R_{2}$ and $C_{1}\setminus\{s\}=C_{2}\setminus\{t\}$ for some $s\in C_{1}$ and $t\in C_{2}$ such that $s\neq t$, or
    • $C_{1}=C_{2}$ and $R_{1}\setminus\{s\}=R_{2}\setminus\{t\}$ for some $s\in R_{1}$ and $t\in R_{2}$ such that $s\neq t$.
@@
[…]
Table 2. Examples of size-$2$ subsets $S$, with formulas for $m^{2}c_{S}(m,n)$ that agree with numeric data. Each minor specification in these examples consists of a contiguous block of rows and a contiguous block of columns, so we can draw them as squares (drawn here on $5\times 5$ matrices).
@@
Conjectures 18 and 21 imply that, if $\lvert{S}\rvert\leq 2$, then the polynomial $m^{\lvert{S}\rvert}c_{S}(m,n)$ is a product of factors that are linear in $m$ and linear in $n$. When we consider subsets $S$ with $\lvert{S}\rvert\geq 3$, we find additional polynomials with this property. In particular, when $S$ has no linked pairs, there seems to be a simple description of $m^{\lvert{S}\rvert}c_{S}(m,n)$ as follows.
@@
Conjecture 22 suggests more generally that the structure of the polynomial $m^{k}c_{S}(m,n)$ is determined by the connected components of linked pairs in $S$ and, moreover, that the contributions of the components are independent of each other. This is the content of Conjecture 24 below. We make this precise by defining the following graph.
@@
Let $S=\begin{smallmatrix}\{\}&\{2\}&\{2,3\}&\{2,4\}\\ \{\}&\{2\}&\{3,4\}&\{3,4\}\end{smallmatrix}$. The first two elements of $S$ form a type-1 link, and the last two form a type-2 link. These are the only links, so $G_{S}$ is the graph on $4$ vertices with two non-adjacent edges. The two connected components are equivalent to the subsets {}{2}{}$\begin{smallmatrix}\{\}&\{2\}\\ \{\}&\{2\}\end{smallmatrix}$ in Example 19 and $\begin{smallmatrix}\{2,3\}&\{3,4\}\\ \{2,3\}&\{2,3\}\end{smallmatrix}$ in Table 2, respectively. The values we computed of $m^{4}c_{S}(m,n)$ are as follows.
@@
This is consistent with $m^{4}c_{S}(m,n)=(m-mn)(n-mn)(2m+n-mn)(2m+3n-mn)$, which is the product of the formulas in Example 19 and Table 2.
@@
Let $S=\begin{smallmatrix}\{2\}&\{3\}&\{2,3\}\\ \{2\}&\{3\}&\{2,3\}\end{smallmatrix}$. Here is the diagram for $S$ on a $5\times 5$ matrix:
The minor specifications ($(\{2\},\{2\})$ and ($(\{3\},\{3\})$ are not linked, but each of the other two pairs is. Therefore $G_{S}$ consists of a single connected component. Here are several values of $m^{3}c_{S}(m,n)$:
@@
Without the value for $6\times 6$ matrices, we cannot interpolate a cubic polynomial in $m$ and $n$. However, by interpolating cubic polynomials in $n$ for the first three rows, we see that $m+n-mn$ is likely a factor. Dividing each value in the table by this factor, we then interpolate a quadratic polynomial to obtain $m^{3}c_{S}(m,n)=(m+n-mn)\left(2m^{2}+2n^{2}+6mn-3m^{2}n-3mn^{2}+m^{2}n^{2}\right)$. The quadratic factor is irreducible.
An obvious question is whether there is a better way to write such polynomials, so that we can see the general structure. In fact there is, using determinants. Since the determinant of a block diagonal matrix is the product of the determinants of the blocks, determinant formulas are good candidates for functions that decompose as products over connected components. In this direction, we next rewrite Conjecture 21 using determinant formulas.
@@
    $m^{2}c_{S}(m,n)=\det\begin{bmatrix}\lvert{R_{1}}\rvert(m+n)-mn&m\\ m&\lvert{R_{2}}\rvert(m+n)-mn\end{bmatrix}.$
@@
    $m^{2}c_{S}(m,n)=\det\begin{bmatrix}\lvert{R_{1}}\rvert(m+n)-mn&n\\ n&\lvert{R_{2}}\rvert(m+n)-mn\end{bmatrix}.$
@@
Along with Conjecture 24, this suggests a possible determinant formula for $m^{\lvert{S}\rvert}c_{S}(m,n)$ for an arbitrary subset $S$. Since the matrix in this determinant formula resembles an adjacency matrix, we introduce the following notation.
@@
For each $S=\begin{smallmatrix}R_{1}&R_{2}&\cdots&R_{k}\vphantom{\{}\\ C_{1}&C_{2}&\cdots&C_{k}\vphantom{\{}\end{smallmatrix}$, define $\operatorname{adj}^{\prime}_{S}(m,n)$ to be the $k\times k$ matrix with the property that, for all $i,j$ satisfying $1\leq i<j\leq k$, the $2\times 2$ submatrix $(\operatorname{adj}^{\prime}_{S}(m,n))_{\{i,j\},\{i,j\}}$ is the matrix in Conjecture 26 for the subset $\begin{smallmatrix}R_{i}&R_{j}\vphantom{\{}\\ C_{i}&C_{j}\vphantom{\{}\end{smallmatrix}$, with the following additional condition. We are building a matrix from a set $S$, and the determinant of this matrix should not depend on the order in which we write the elements of $S$. Three of the cases in Conjecture 26 are symmetric in $(R_{i},C_{i})$ and $(R_{j},C_{j})$, but the first case is not. Therefore, if $(R_{i},C_{i})$ and $(R_{j},C_{j})$ form a type-1 link with $\lvert{R_{i}}\rvert=\lvert{R_{j}}\rvert+1$, we permute the rows and columns relative to the first case in Conjecture 26, so that $(\operatorname{adj}^{\prime}_{S}(m,n))_{\{i,j\},\{i,j\}}$ is
$\begin{bmatrix}\lvert{R_{i}}\rvert(m+n)-mn&-n\\ m&\lvert{R_{j}}\rvert(m+n)-mn\end{bmatrix}.$
@@
$\operatorname{adj}^{\prime}_{S}(m,n)=\begin{bmatrix}m+n-mn&0&m\\ 0&m+n-mn&m\\ -n&-n&2m+2n-mn\end{bmatrix}.$
Indeed, $\det\operatorname{adj}^{\prime}_{S}(m,n)$ is equivalent to the formula in Example 25 for $m^{3}c_{S}(m,n)$.
In defining $\operatorname{adj}^{\prime}_{S}(m,n)$, we inadvertently made choices about the signs of the off-diagonal entries. Namely, negating the off-diagonal entries of the $2\times 2$ matrices in Conjecture 26 doesn’t change their determinants, but of course negating a pair of off-diagonal entries in a larger matrix can change its determinant. Therefore there is no reason to expect $\det\operatorname{adj}^{\prime}_{S}(m,n)$ to give the intended coefficient for all subsets $S$, and in fact it does not.
@@
Let $S=\begin{smallmatrix}\{2\}&\{2\}&\{2,3\}\\ \{2\}&\{3\}&\{2,3\}\end{smallmatrix}$. The minor specifications ($(\{2\},\{2\})$ and ($(\{2\},\{3\})$ form a type-2 link, and each forms a type-1 link with ($(\{2,3\},\{2,3\})$. We have
$\operatorname{adj}^{\prime}_{S}(m,n)=\begin{bmatrix}m+n-mn&m&m\\ m&m+n-mn&m\\ -n&-n&2m+2n-mn\end{bmatrix}.$
The formula $\det\operatorname{adj}^{\prime}_{S}(m,n)$ does not produce the following values we computed for $m^{3}c_{S}(m,n)$.
$\begin{array}[]{r|rrrr}&n=3&4&5&6\\ \hline\cr m=3&0&-16&-80&-216\\ 4&-6&-128&-490&-1200\\ 5&-36&-432&-1500&-3528\\ 6&-108&-1024&-3380\end{array}$
However, if we alter the signs of the off-diagonal entries, we can get a determinant formula that produces these values, namely
$m^{3}c_{S}(m,n)=\det\begin{bmatrix}m+n-mn&m&m\\ m&m+n-mn&-m\\ -n&n&2m+2n-mn\end{bmatrix}.$
Further examples suggest that the signs of the off-diagonal entries in $\operatorname{adj}^{\prime}_{S}(m,n)$ can always be altered in such a way that its determinant gives the value of $m^{\lvert{S}\rvert}c_{S}(m,n)$. However, the graph $G_{S}$ does not determine the signs, as the following example shows.
@@
For the subset $S=\begin{smallmatrix}\{2,3,4\}&\{2,3,4\}&\{2,3,4\}&\{2,3,4\}&\{2,3,4\}\\ \{2,3,4\}&\{2,3,5\}&\{2,5,7\}&\{2,6,7\}&\{2,4,6\}\end{smallmatrix}$, the graph $G_{S}$ is a $5$-cycle. The determinant of
$\begin{bmatrix}3m+3n-mn&m&0&0&-m\\ m&3m+3n-mn&-m&0&0\\ 0&-m&3m+3n-mn&m&0\\ 0&0&m&3m+3n-mn&-m\\ -m&0&0&-m&3m+3n-mn\end{bmatrix}$
gives the values of $m^{5}c_{S}(m,n)$. For the subset $T=\begin{smallmatrix}\{2,3,4\}&\{2,3,4\}&\{2,3,4\}&\{2,3,4\}&\{2,3,4\}\\ \{2,3,4\}&\{2,3,5\}&\{3,5,6\}&\{4,5,6\}&\{2,4,6\}\end{smallmatrix}$, the graph $G_{T}$ is also a $5$-cycle, but we need to alter one pair of signs; the determinant of
$\begin{bmatrix}3m+3n-mn&m&0&0&-m\\ m&3m+3n-mn&+m&0&0\\ 0&+m&3m+3n-mn&m&0\\ 0&0&m&3m+3n-mn&-m\\ -m&0&0&-m&3m+3n-mn\end{bmatrix}$
gives the values of $m^{5}c_{T}(m,n)$.
To determine the signs of the off-diagonal entries in general, we used ChatGPT o3-mini-high [18]. We presented the entire framework of the question, including our conjectures about the determinant formula for the coefficients, along with the previous example of two subsets with the same graph that require two different sign assignments. [^3] For the conversation, see https://chatgpt.com/share/67aba51d-c0f0-800d-a9cb-d3533aadc17c. ChatGPT’s responses were self-contradictory (it claimed that the product around cycles must be positive) and not entirely prescriptive, but it suggested that, for a linked pair of minor specifications $(R_{i},C_{i})$ and $(R_{j},C_{j})$, the relative positions of the row and column indices are relevant. Namely, the signs of the corresponding off-diagonal entries should reflect the permutations mapping $R_{i}$ to $R_{j}$ and $C_{i}$ to $C_{j}$ as ordered lists. We hadn’t considered this ourselves since the row and column orders play no role up to this point. This was enough direction for us to find rules for the signs that agreed with all numeric data we had computed. For a type-1 link, the index sets have different sizes, so to obtain a permutation we append the missing element to the smaller list. For a type-2 link, the index sets have the same sizes but can differ in two elements, so to obtain a permutation we require that those two elements get mapped to each other.